

********************************************************************
UARS WEEKLY STATUS REPORT - 10 SEP 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data this past
week without apparent problems.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM/AXIS activated between September 4 and September 9. While AXIS
was active, it produced energy deposition from 16 pixels based on precipitating
electrons derived from measured X-ray fluxes.
The PEM spectrometers continue to be active under PMON control. While
under PMON control, the PEM instrument computed energy deposition from its
particle spectrometers.
The FOT and the PEM team are updating the PEM activation sequence which
effects the PEM/HEPS (> 30 keV particles). Current activation sequence
reduces the activation time by about 1 minute. Additional analization shows
that the activation sequence may be reduced by about another minute. PEM
PMON activation is set for the beginning of power orbit day. The beginning of
power orbit day is defined as when:
array current greater than 5 amps, and
array voltage greater than 53.5 volts, and
any battery current greater than 2.0 amps
The end of power orbit day (PEM deactivation) is defined as when:
time in power orbit day greater than 4 minutes, and
array current less than 9 amps, and
all battery currents less than 0.2 amps
Power orbit day periods vary depending upon beta angle:
about 39-40 minutes at yaw
about 50-51 minutes at max beta
Current activation takes about 7.5 min for full activation. During the
activation sequence, various components of PEM become active at all different
times, beginning with the PEM magnetometer (at about on activation time) and
ending with the HEPS-2 components (7.5 min later).
***********
HALOE
***********
HALOE continues successfully operations with all engineering parameters
with nominal operating range. A HALOE data validation meeting was held
this week to review the data from May, June and July 1999. Results
continue to be excellent.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII was activated on September 9, and is operating normally.
Operations will continue until September 14.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
14 days CLA (7.0) PROD (L1-L3) (473R-238C-14F=221L)
0 days GPI (4.05) PROD (CLAES L3B) (191R-164C-2F=25L)
0 days WIN (5.11) PROD (L1-L3) (1540R-1028C-33F=479L)
The CDHF continued integrating the "UARS" environment on CDHFA/B.
CDHF pending software upgrades: none
***********
FOT
***********
This report covers 28 August, 1999 (Orbit 43506, GMT Day 240) through
3 September, 1999 (Orbit 43611, GMT Day 246).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 239/0007 43492 Off
The beta angle ranged between the angles of 26.9 degrees to 49.5 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
51 degrees on 07 September 1999 (DOY 250).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.7 to 1.0 degrees C. Additional battery performance data for each
battery on 27 August 1999 and 03 September 1999 are provided in Appendix D.
The clock error ranged from 11.1 to -10.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 03
% WINDII = 01
% PEM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
One (3) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/242 - 2287.496953
% Transponder B: 99/243 - 2287.499030
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/246 show the observatory data loss to be 24 hours, 55
minutes, 09.428 seconds (an increase of 18.912 seconds since the last report
period). This is a 0.0358 percent data loss which equals a 99.9642 percent
data capture for the mission.
The increase of 0.224 seconds occurred over one (1) days:
% 0.224 seconds on DOY 241 due to data hits
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to investigate the feasibility of a new PMON action to change
battery V/T curves based on loading to mitigate battery risks in sun pointing
safehold.
PEMUs timing gap was adjusted. The power turn on sequence was increased by 48
seconds.
YEAR 2000 OPERABILITY SUPPORT
Ephemeris products were given to Charles Coleman for him to update the Y2K
scenario for testing next week.
MEETINGS
None to report this week.
FUTURE SIGNIFICANT EVENTS
September, 1999: A meeting is planned for the beginning of September to
discuss the UARS Safehold configuration and concerns for the MPS and its
batteries. This meeting will be a closed meeting with only the FOT and
specialized engineering support attending.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 3 September 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
43546 242/1436 99-308 OBC Clk Adj fr 0.00 to 20.6 msec/day
43572 244/0937 99-309 OBC Clk Adj fr 20.6 to 0.00 msec/day
43579 244/1928 99-310 PEM PMON/PEM
43596 245/2244 99-311 WINDII Power OFF
43603 246/0954 99-312 OBC Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 3 September 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
No Out-of-Limits to report this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 3 September 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ------- -----------------
43546 242/14:40 (30 Aug 99) from 00.0 to 20.6 msec/day
43572 244/09:41 (01 Sept 99) from 20.6 to 00.0 msec/day
43603 246/09:56 (03 Sept 99) from 00.0 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 3 September 1999
27 August 1999 - GMT Day 239 Beta = 26.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- --------- ---------- ----- ------- ------
1 -112.0/-134.4 -2.73/-4.31 +0.0/+0.0 16.5 0.000 0.0
2 +67.2/+0.0 +6.44/+4.51 +30.8/-15.2 24.3 1.023 22.0
3 +39.2/-5.6 +5.47/+3.87 +30.4/-16.0 24.2 1.022 21.8
03 Sept 1999 - GMT Day 232 Beta = 49.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ----- ------- ------
1 -112.0/-134.4 -3.05/-4.31 +0.0/+0.0 16.5 0.000 0.0
2 +39.2/+0.0 +5.47/+4.51 +26.4/-15.2 25.3 1.019 21.4
3 +33.6/-11.2 +4.51/+3.87 +25.6/-15.2 25.1 1.023 21.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 3 September 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-044 FOT 43543 242/1114 N/A AP Crash - NIU down
99-045 FOT 43574 244/1236 N/A AP Crashed-switched Pack
99-046 FOT 43576 244/1607 N/A TAC 4 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-035 FOT 43442 235/1554 N/A TAC 6 Crashed
99-036 FOT 43445 235/2212 N/A TAC 6 Crashed
99-037 FOT 43458 236/1855 21265 Late Acq / sched.conflict
99-038 FOT 43442 235/1554 N/A TAC 6 Crashed
99-039 FOT see list Aug 99 14170 Generic Late Acq.
99-041 FOT 43445 235/2212 N/A TAC 6 Crashed
99-042 FOT 43458 236/1855 21265 Late Acq / sched.conflict
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-045 FOT 43574 244/1236 N/A AP Crashed-switched Pack
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-033 FOT 42856 196/1323 N/A TMON #1 trip
99-040 FOT 43187 218/1634 21201 NASCOM problem at WSGT
99-043 FOT 43458 236/1855 21265 Late Acq / sched.conflict
SEVNTFW99/36